five 10- resistor are connected in parallel and the combination is connected to 10-V emf device. what is the current device


1
Expert's answer
2022-05-06T15:34:21-0400
"R_{net}=\frac{R}{10}=\frac{10\:\Omega}{10}=1\: \Omega"

"I=\frac{V}{R_{net}}=\rm\frac{10\: V}{1\; \Omega}=10\: A"